Hoodmemes:

The area of a rectangular field is equal to 300 square meters. Its perimeter is equal to 70 meters. Find the length and width of this rectangle.

cpl106:

Let the length and the width of the field be L and W Perimeter = 2(W + L) = 70 => W + L = 35 => W = 35 - L Area = W * L = 300 => (35 - L)* L = 300 => 35 L - L^2 = 300 => L^2 - 35 L + 300 = 0 => L^2 - 15 L - 20 L + 300 = 0 => L( L-15) - 20( L- 15) = 0 => (L - 20) ( L -15) = 0 So L can be 20 and 15 W = 15 and 20 As length is greater than the width, we have the length equal to 20 and the width equal to 15.

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @cpl106 Let the length and the width of the field be L and W Perimeter = 2(W + L) = 70 => W + L = 35 => W = 35 - L Area = W * L = 300 => (35 - L)* L = 300 => 35 L - L^2 = 300 => L^2 - 35 L + 300 = 0 => L^2 - 15 L - 20 L + 300 = 0 => L( L-15) - 20( L- 15) = 0 => (L - 20) ( L -15) = 0 So L can be 20 and 15 W = 15 and 20 As length is greater than the width, we have the length equal to 20 and the width equal to 15. \(\color{#0cbb34}{\text{End of Quote}}\) thanks,but u didnt need to give me the direct answer

cpl106:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Hoodmemes \(\color{#0cbb34}{\text{Originally Posted by}}\) @cpl106 Let the length and the width of the field be L and W Perimeter = 2(W + L) = 70 => W + L = 35 => W = 35 - L Area = W * L = 300 => (35 - L)* L = 300 => 35 L - L^2 = 300 => L^2 - 35 L + 300 = 0 => L^2 - 15 L - 20 L + 300 = 0 => L( L-15) - 20( L- 15) = 0 => (L - 20) ( L -15) = 0 So L can be 20 and 15 W = 15 and 20 As length is greater than the width, we have the length equal to 20 and the width equal to 15. \(\color{#0cbb34}{\text{End of Quote}}\) thanks,but u didnt need to give me the direct answer \(\color{#0cbb34}{\text{End of Quote}}\) oh my bad

cooksd2023:

Let the length and the width of the field be L and W Perimeter = 2(W + L) = 70 i was to late but i think we were all on the same page

Jade123:

The person who makes it has no need of it; the person who buys it has no use for it. The person who uses it can neither see nor feel it. What is it?

cooksd2023:

idk Shakespeare

Jade123:

no try again

Jade123:

cpl you wanna try?

cooksd2023:

coffin.

Jade123:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @cooksd2023 coffin. \(\color{#0cbb34}{\text{End of Quote}}\) correct

cooksd2023:

wow

Jade123:

I turn once, what is out will not get in. I turn again, what is in will not get out. What am I?

cooksd2023:

a door

Jade123:

nope but close

cooksd2023:

umm a key

Jade123:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @cooksd2023 umm a key \(\color{#0cbb34}{\text{End of Quote}}\) correct

cooksd2023:

omg im smart

Jade123:

two in a corner, one in a room, zero in a house, but one in a shelter. What is it?

cooksd2023:

that one is a lil hard umm

Jade123:

its a letter

cooksd2023:

the letter r ?

Jade123:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @cooksd2023 the letter r ? \(\color{#0cbb34}{\text{End of Quote}}\) correct

cooksd2023:

wow i thought i got it wrong i wish i can get my test right like these riddles

Jade123:

I am a word that begins with the letter “i.” If you add the letter “a” to me, I become a new word with a different meaning, but that sounds exactly the same. What word am I?

Jade123:

lol

cooksd2023:

a